how do you determine the exact value of tan2x if cotx=4/3,0
Answer by lwsshak3(11628)

determine the exact value of tan2x if cotx=4/3,0 cotx=1/tanx=4/3
tanx=3/4

..
check:
tanx=3/4
x=36.87
2x=73.74
by calculator:
tan2x=tan 73.74˚≈3.4286...
exact value as computed=48/24≈3.4285...